OpenStudy (anonymous):

integrate from 9 to 1 (sqrt(u) - 2u^2)/ u DU not sure what to substitute in or how to start

OpenStudy (freckles):

separate the fraction

OpenStudy (dumbcow):

\[\int\limits \frac{\sqrt{u} -2u^2}{u} = \int\limits u^{-1/2} - 2u\]

OpenStudy (solomonzelman):

came across... where are your du's ?